# SMC-X: A Distributed Scalable Monte Carlo Simulation Method for Chemically Complex Alloys

### v1

@@ 3.1 Performance
Figure 4: Comparison of different ML accelerated atomistic simulation methods in terms of atom per chip and throughput per chip.
@@
LAMMPS 27 Allegro HIV 5,120 Perlmutter 100 M $1.27\times 10^{10}$ $2.47\times 10^{6}$}
LAMMPS 47 DP H2O 90,000 Sunway 29 B $1.14\times 10^{10}$} $1.26\times 10^{5}$
LAMMPS 48 TensorMD W 32,000 Sunway 51.8 B $\mathbf{{1.30\times 10^{10}}}$ $4.06\times 10^{5}$
GPUMD 49 UNEP-v1 Cu 8 A100 1 node 100 M $1.50\times 10^{8}$} $1.86\times 10^{7}$
GPUMD 50 NEP89 HEA 1 H800 1 node 15 M $7.0\times 10^{6}$} $7.0\times 10^{6}$
@@
The performance is measured with three metrics: number of atoms, which measure the spatial scale; simulation throughput, which measures time to solution; and scaling efficiency, which measures the scale-up potential with increasing computational resources. As shown in Fig. 4, the EPI model exhibits exceptional computational efficiency on a single GPU, achieving a throughput of $3.5\times 10^{9}$ $\text{atom}\cdot\text{step}/\text{s}$—two orders of magnitude faster than the qSRO_GPU result of $3.7\times 10^{7}$ $\text{atom}\cdot\text{step}/\text{s}$. In comparison, the single-NUMA performance of qSRO_CPU reaches $3.5\times 10^{6}$ $\text{atom}\cdot\text{step}/\text{s}$, or $2.5\times 10^{7}$ $\text{atom}\cdot\text{step}/\text{s}$ per CPU. The throughput of SMC-X not only far exceeds that of traditional Monte Carlo methods but also outperforms state-of-the-art molecular dynamics implementations, as shown in Fig. 4 and summarized in Tab. 2. The advantage is even more pronounced in terms of the number of atoms supported on a single chip: the SMC-GPU implementation accommodates up to 1 billion atoms on a single GPU, whereas the best reported MD simulation to date reaches only 15 million atoms per GPU, as shown in the row of NEP89 in Tab. 2. As mentioned, these performance gains stem from key implementation optimizations, including on-the-fly computation and tight coupling between the atomistic simulation and machine learning models, which collectively minimize data movement and maximize computational efficiency.
@@ 3.2 Evolution of chemical defects
The Warren–Cowley short-range order (SRO) parameter 51, 31, denoted as $\alpha_{m}^{pp^{\prime}}$, is widely used to quantify local chemical ordering. It is defined as
@@
where $c_{p}$ is the concentration of species $p$, and $P^{p|p^{\prime}}_{m}$ is the conditional probability of finding an atom of type $p$ in the $m$-th neighbor shell around a central atom of type $p^{\prime}$. The SRO parameters are obtained by averaging over the Monte Carlo configurations. By definition, $\alpha=0$ corresponds to perfect randomness, $\alpha<0$ indicates a tendency toward ordering between $p$ and $p^{\prime}$, and $\alpha>0$ reflects a tendency toward segregation. Figure 6 shows the nearest-neighbor (NN) SRO parameters obtained from Monte Carlo simulations of a one-million-atom FeCoNiAlTi system. At high temperature, the alloy remains nearly random, whereas at lower temperatures distinct SRO develops. In particular, Ti and Ni atoms display a strong tendency to segregate from Fe and Co, with the Ti–Fe pair showing the strongest separation. This behavior is consistent with previous experimental 7 and theoretical 22 studies, which reported the formation of $L1_{2}$-type nanoprecipitates enriched in Ni and Ti within a Fe- and Co-rich matrix.
Figure 6: The short range order parameters calculated with the 1 million atom supercell. (a) The SROs for different elements pairs; (b) The SROs for the same element pairs.
To investigate microstructure evolution, we performed large-scale Monte Carlo simulations using the EPI model, with a supercell comprising 1,066,867,200 atoms (dimensions: $672\times 630\times 630\times 4$). The computing platform is the two node, 16 H800 GPU cluster, and the ML model is the EPI model to maximize the simulation speed. We decrease the temperature from 2000 K to 1000 K, with tens of thousands of MC steps at each temperature, except for a temperature we refer to as the “aging” temperature $T_{aging}$, at which the simulation run for 3 million MC steps . The X-Y face of the system for $T_{aging}=1000$ K experiment is shown in Fig. 7 (a, b), where we used 8 GPUs for lattice parallelism and 2 temperature degree of parallelism, with $\Delta T=0.06$ K. As mentioned in the method section, such a small $\Delta T$ is for reasonable acceptance ratio of RE in large supercells. The results for $T_{aging}=1500$ K is shown in Fig. 7 (c, d), where no RE is used, and the lattice is parallelized over the total 16 GPUs. From these figures, it can be seen that the NPs rich in Ni and Ti already forms at 1500 K. Upon further cooling, nanoprecipitates gradually grow in size. Moreover, the comparison between Fig. 7 (a) and (c) suggests that the two temperature replicas demonstrate very little effect, which is consistent with our expectation, due to the tiny value of $\Delta T$. Finally, it can be seen that the size of NPs increases as $T_{aging}$ switched from 1000 K to 1500 K, indicating the high rate of NP formation at elevated temperature.
@@ 3.3 APT sample
Figure 8: Comparison of APT sample from the simulation with previous experimental and theoretic studies. (a) The APT sample extracted from SMC-X simulation in Ref. 22. (b) The APT reconstructed sample image of the same materials from experiment 7, where MCM represents multicomponent matrix, and MCINP represents multicomponent intermetallic nanoparticle. (c) The APT sample from this work, using 3 million MC steps at 1000 K. (d) The APT sample extracted from Fig. 7 d, using 3 million MC steps at 1500 K. (a-c) blue represents MCINP, and red represents MCM.
To reveal further details of the NPs, we extract an atom-probe-tomography (APT) sample from the simulation results and compared them with previous experimental and theoretical studies, as shown in Fig. 8. First, we see that the previous study shown in Fig. 8(a) clearly underestimates the sizes of the NPs, which is about 7 nm in diameter, as compared to the experimental result in Fig. 8 (b), which is about 30 nm. Second, we see that this gap between experiment and theory can be reduced by increasing the MC steps from 0.1 million to 3 million, which enhances the size of the NP to about 14 nm. Finally, note that the experimental sample is synthesized by homogenizing at 1423 K for 2 h and aged at 1053 K for 4 h. The 3 million steps at 1500 K in Fig. 8(d) can be seen as an effort to mimic the homogenization effect. It can be seen that the size of the NP can be further increased to about 20 nm. These observations reveal that larger number of steps are needed for the system to reach equilibrium, and the discrepancy in the size of the nanoparticles in Ref. 7 and Ref. 22 can be narrowed by increasing the simulation steps to mimic the whole heat-treatment process. We emphasize that the distributed implementation presented in this is the reason that we can run such a large-number of MC steps in a time budget of a few days.
Although there is no time dependence in MC method, the time scale of the simulation can be estimated using a simple diffusion model, similar to kinetic MC. A well-known model to describe the diffusion-controlled coarsening is Lifshitz–Slyozov–Wagner (LSW) theory, which gives
@@
Of course, this is a very crude estimation, but it indicate that our MC simulation indeed reaches the time regime of real physical process of heat-treatment.
@@ 4 Conclusions
In this work, we have presented a distributed framework for the SMC-X method, demonstrating its versatility with implementations on both CPU and GPU architectures. By leveraging hierarchical parallelism, data-locality optimization, and computation–communication overlap, the distributed SMC-X achieves exceptional performance in system size, simulation throughput, and scaling efficiency. When combined with DFT-trained machine learning models, SMC-X attains first-principles accuracy while extending both the spatial and temporal scales of atomistic simulations to previously inaccessible regimes. Such capability is essential for predicting large-scale chemical defects in chemically complex alloys and, in turn, for elucidating the origin of their exceptional properties. This claim is exemplified by our billion-atom simulations of FeCoNiAlTi, where the extended spatial and temporal scales prove critical for bridging the gap between experimental observations and theoretical predictions. Altogether, our results demonstrate that the SMC-X + ML framework sheds light on the challenging yet critical problem of simulation-driven design of chemically complex materials.
@@
All the dadta data will be publically available in Zenodo upon the acceptance of this paper.
@@
The work of X. Liu and F. Zhou was supported by the National Natural Science Foundation of China under Grant 12404283. X. Liu acknowledges helpful discussion with Zongrui Pei on precipitate sizes in alloy.
@@
The authors declare no competing interests.

### v2

@@ 3.1 Performance
(a) Atoms per chip comparison
(b) Throughput per chip comparison
(c) Overall performance metrics
Figure 4: Comparison of different ML accelerated atomistic simulation methods: (a) shows the number of atoms that can be simulated per chip, (b) displays the computational throughput achievable per chip, and (c) comparison of the performance of our results with previous SOTA in terms of total $N_{atom}$ ($x$-axis), total throughput ($y$-axis), throughput per chip (color), and the number of chips (circles with radius proportional to $N_{Chip}^{1/3}$). See original data at Tab. 2.
@@
LAMMPS 10.1145/3581784.3627041 Allegro HIV 5,120 A100 Perlmutter 100 M $1.27\times 10^{10}$ $2.47\times 10^{6}$}
LAMMPS 10880101 DP H2O 90,000 CPU Sunway 29 B $1.14\times 10^{10}$} $1.26\times 10^{5}$
LAMMPS 10.1145/3503221.3508425 DP H2O 27,360 V100 Summit 3.9B $1.12\times 10^{10}$ $4.11\times 10^{5}$
LAMMPS 10.1145/3503221.3508425 DP H2O 9,936 CPU Fugaku 1.5B $2.17\times 10^{9}$ $2.18\times 10^{5}$
LAMMPS 10.1145/3458817.3487400 SNAP C 27,900 V100 Summit 20B $\mathbf{2.89\times 10^{10}}$ $1.08\times 10^{6}$
LAMMPS 10.1145/3710848.3710882 TensorMD W 32,000 CPU Sunway 51.8 B $1.30\times 10^{10}$}} $4.06\times 10^{5}$
GPUMD MLPAlloys2024 UNEP-v1 Cu 8 A100 1 node 100 M $1.50\times 10^{8}$} $1.86\times 10^{7}$
GPUMD liang2025nep89universalneuroevolutionpotential NEP89 HEA 1 H800 1 node 15 M $7.0\times 10^{6}$} $7.0\times 10^{6}$
@@
Performance is measured with three metrics: number of atoms, which measure spatial scale; simulation throughput, which measures time to solution; and scaling efficiency, which measures the scaling potential with increasing computational resources. As shown in Fig. 6(b), the EPI model exhibits exceptional computational efficiency on a single GPU, achieving a throughput of $3.5\times 10^{9}$ $\text{atom}\cdot\text{step}/\text{s}$—two orders of magnitude faster than the qSRO_GPU result of $3.7\times 10^{7}$ $\text{atom}\cdot\text{step}/\text{s}$. In comparison, the single-NUMA performance of qSRO_CPU reaches $3.5\times 10^{6}$ $\text{atom}\cdot\text{step}/\text{s}$, or $2.5\times 10^{7}$ $\text{atom}\cdot\text{step}/\text{s}$ per CPU. The throughput of SMC-X not only far exceeds that of traditional Monte Carlo methods, but also outperforms state-of-the-art molecular dynamics implementations, as shown in Fig. 4 and summarized in Tab. 2. It is noteworthy that SMC-X, utilizing merely 32 H800 GPUs, can simulate systems of 128 billion atoms, which is a scale that surpasses all other ML-accelerated MC or MD simulations, including those performed on some of the world’s most powerful supercomputers, as shown in Fig. 4(c). The advantage is even more pronounced in terms of the number of atoms supported on a single chip: the SMC-GPU implementation accommodates up to 4 billion atoms on a single GPU by running 4 MPI ranks simultaneously, whereas the best reported MD simulation to date reaches only 15 million atoms per GPU, as shown in the row of NEP89 in Tab. 2. As mentioned, these performance gains stem from key implementation optimizations, including on-the-fly computation and tight coupling between atomistic simulation and machine learning models, which collectively minimize data movement and maximize computational efficiency.
@@
Finally, we comment on the limitations of current SMC-X implementations and its future potential. The present implementation requires kernel-level programming to optimize the machine-learning (ML) potential, which enables peak performance but makes the integration of additional potentials labor-intensive. Consequently, we have so far implemented only a simple qSRO model as a representative of general short-range models. This naturally raises the question of whether such performance can be sustained with more complex ML interatomic potentials (MLIPs). To investigate this, we performed a roofline analysis of the SMC-GPU code using NVIDIA Nsight tools. The results show that the current implementation is bounded primarily by integer arithmetic for on-the-fly index calculations and by HBM bandwidth for data movement, whereas the floating-point units, including CUDA vector units and Tensor Cores, remain significantly underutilized due to the low theoretical arithmetic intensity of about 1. Given that modern GPUs are designed for workloads with arithmetic intensities in the hundreds or higher, it is reasonable to expect that incorporating more complex models with higher computational intensity will not compromise performance. Indeed, in a prototype implementation using a three-layer multilayer perceptron (MLP) with a hidden dimension of 256, we observe that the excellent speed reported for qSRO is retained. Looking ahead, accommodating fully pretrained MLIPs may require a modest redesign of the SMC-X code architecture, potentially introducing a small performance overhead. Nevertheless, the method would remain exceptionally fast compared with conventional Monte Carlo approaches, as the dominant speedup arises from the SMC-X algorithm itself. These enhancements will be explored in future work.
@@ 3.2 Evolution of chemical defects
The Warren–Cowley short-range order (SRO) parameter PhysRev.77.669, LIU2021110135, denoted $\alpha_{m}^{pp^{\prime}}$, is widely used to quantify local chemical ordering. It is defined as
@@
where $c_{p}$ is the concentration of species $p$, and $P^{p|p^{\prime}}_{m}$ is the conditional probability of finding an atom of type $p$ in the $m$-th neighbor shell around a central atom of type $p^{\prime}$. The SRO parameters are obtained by averaging over the Monte Carlo configurations. By definition, $\alpha=0$ corresponds to perfect randomness, $\alpha<0$ indicates a tendency toward ordering between $p$ and $p^{\prime}$, and $\alpha>0$ reflects a tendency toward segregation. Figure 6 shows the nearest-neighbor (NN) SRO parameters obtained from Monte Carlo simulations of a one-million-atom FeCoNiAlTi system. At high temperature, the alloy remains nearly random, whereas at lower temperatures distinct SRO develops. In particular, Ti and Ni atoms display a strong tendency to segregate from Fe and Co, with the Ti–Fe pair showing the strongest separation. This behavior is consistent with previous experimental Yang933 and theoretical liuNPJ2025 studies, which reported the formation of $L1_{2}$-type nanoprecipitates enriched in Ni and Ti within a Fe- and Co-rich matrix.
(a) SRO of different element pairs
(b) SRO of identical element pairs
Figure 6: The chemical short-range order parameters (SROs) calculated with the 1 million atom supercell, estimated with 9 configuration samples. The error bars are estimated with the difference between the largest and smallest ones among the 9 SRO values.(a) The SROs for different elements pairs; (b) The SROs for the same element pairs.
To investigate microstructure evolution, we performed large-scale Monte Carlo simulations using the EPI model, with a supercell comprising 1,066,867,200 atoms (dimensions: $672\times 630\times 630\times 4$). The computing platform is the two-node, 16-H800 GPU cluster, and the ML model is the EPI model to maximize the simulation speed. We decrease the temperature from 2000 K to 1000 K, with tens of thousands of MC steps at each temperature, except for a temperature we refer to as the “aging” temperature $T_{aging}$, at which the simulation runs for 3 million MC steps . The X-Y face of the system for $T_{aging}=1000$ K experiment is shown in Fig. 7 (a, b), where we used 8 GPUs for lattice parallelism and 2 temperature degrees of parallelism, with $\Delta T=0.06$ K. As mentioned in the Methods section, such a small $\Delta T$ is for a reasonable acceptance ratio of RE in large supercells. The results for $T_{aging}=1500$ K is shown in Fig. 7 (c, d), where no RE is used, and the lattice is parallelized over the total 16 GPUs. From these figures, it can be seen that the NPs rich in Ni and Ti already form at 1500 K. Upon further cooling, nanoprecipitates gradually grow in size. Moreover, the comparison between Fig. 7 (a) and (c) suggests that the two temperature replicas demonstrate very little effect, which is consistent with our expectation, due to the tiny value of $\Delta T$. Finally, it can be seen that the size of NPs increases as $T_{aging}$ switched from 1000 K to 1500 K, indicating the high rate of NP formation at elevated temperature.
@@ 3.3 APT sample
Figure 8: Comparison of APT sample from the simulation with previous experimental and theoretic studies. (a) The APT sample extracted from SMC-X simulation in Ref. liuNPJ2025. Reproduced from Ref. liuNPJ2025. Available under a $C$-BY NC-ND 4.0 license. Copyright Xianglin Liu et al. (b) The APT reconstructed sample image of the same materials from experiment Yang933, where MCM represents multicomponent matrix, and MCINP represents multicomponent intermetallic nanoparticle. Reproduced with permission from Ref. Yang933. Copyright 2018 AAAS. (c) The APT sample from this work, using 3 million MC steps at 1000 K. (d) The APT sample extracted from Fig. 7 d, using 3 million MC steps at 1500 K. (a-c) blue represents MCINP, and red represents MCM. (e) Averaged compositions obtained from 9 nearby cylinderical specimen of r=12ar=12\;a and l=72al=72\;a, where $a$ is the lattice constant.
To reveal further details of the NPs, we extract an atom-probe-tomography (APT) sample from the simulation results and compare them with previous experimental and theoretical studies, as shown in Fig. 8. First, we see that the previous study shown in Fig. 8(a) clearly underestimates the sizes of the NPs, which is about 7 nm in diameter, compared to the experimental result in Fig. 8 (b), which is about 30 nm. Second, we see that this gap between experiment and theory can be reduced by increasing the MC steps from 0.1 million to 3 million, which increases the size of the NP to about 14 nm. Finally, note that the experimental sample is synthesized by homogenizing at 1423 K for 2 h and aged at 1053 K for 4 h. The 3 million steps at 1500 K in Fig. 8(d) can be seen as an effort to mimic the homogenization effect. It can be seen that the size of the NP can be further increased to about 20 nm. These observations reveal that a larger number of steps are needed for the system to reach equilibrium, and the discrepancy in the size of the nanoparticles in Ref. Yang933 and Ref. liuNPJ2025 can be narrowed by increasing the simulation steps to mimic the whole heat-treatment process. We emphasize that the distributed implementation presented in this is the reason that we can run such a large-number of MC steps in a time budget of a few days. The chemical composition at 1000 K is analyzed, as shown in Fig 8 (e). The composition profile is extracted from 9 nearby cylindrical specimen of radius r=12ar=12\;a and length of l=72al=72\;a, where $a$ is the lattice constant. It can be seen that Fe and Co are favored in the disordered matrix phase, while Ni and Ti and strongly favored in the ordered (Ni,Co)3-(Ti,Al) $L1_{2}$ phase. These results agrees well with both experiments Yang933 and previous theoretical studies liuNPJ2025, WANG2025120635. The error bar is the standard deviation of the compositions.
Although there is no time dependence in the MC method, the time scale of the simulation can be estimated using a simple diffusion model, similar to kinetic MC. A well-known model to describe the diffusion-controlled coarsening is Lifshitz–Slyozov–Wagner (LSW) theory, which gives
@@
Of course, this is a very crude estimation, but it indicates that our MC simulation indeed reaches the time regime of a real physical process of heat-treatment.
@@ 4 Conclusions
In this work, we have presented a distributed framework for the SMC-X method, demonstrating its versatility with implementations on both CPU and GPU architectures. Using hierarchical parallelism, data-locality optimization, and computation–communication overlap, the distributed SMC-X achieves exceptional performance in simulation throughput and scaling efficiency. In particular, it achieves the largest system size for ML-accelerated atomistic simulation, reaching 128 billion atoms using only 32 H800 GPUs, a much smaller hardware budget compared to previous results obtained with the largest supercomputers. When combined with DFT-trained machine learning models, SMC-X attains first-principles accuracy while extending both the spatial and temporal scales of atomistic simulations to previously inaccessible regimes. Such capability is essential for predicting large-scale chemical defects in chemically complex alloys and, in turn, for elucidating the origin of their exceptional properties. This claim is exemplified by our billion-atom simulations of FeCoNiAlTi, where extended spatial and temporal scales are critical for bridging the gap between experimental observations and theoretical predictions. Together, our results demonstrate that the SMC-X + ML framework sheds light on the challenging but critical problem of simulation-driven design of chemically complex materials.
@@
[…]
@@
The work of X. Liu and F. Zhou was supported by the National Natural Science Foundation of China under Grant 12404283. The work of F. Zhou was also supported by the High-Level Talent Start-up Fund provided by Xiangnan University. X. Liu acknowledges helpful discussion with Zongrui Pei on precipitate sizes in alloy.
@@
The authors declare no conflict of interest.